OpenStudy (daniel.ohearn1):

What is the integral evaluated from zero to pi/4 of (sin^4 of x * cos^2 of x ) dx ?

OpenStudy (tkhunny):

Please demonstrate your efforts.

OpenStudy (danjs):

page 2 has the trig sub thigns to remind of some tricks

OpenStudy (solomonzelman):

\(\color{#0000a0 }{ \displaystyle \int\limits_0^{\pi/4} \left(\sin^4x\cdot \cos^2x\right){\tiny~}{\mathrm d}x}\) This?

OpenStudy (solomonzelman):

YOu can write the integrand as cosines of 2x, 4x, and 6x.

OpenStudy (tkhunny):

I'm not sure you have to go to ALL that trouble. This looks promising: \(\dfrac{1}{2}\int\left(\sin^{3}{x}\cdot\cos(x)\right)\cdot\sin(2x)\;dx\)